# Adjusted Connections I: Differential Cocycles for Principal Groupoid Bundles with Connection

### v1

@@ 6.3 General statements on existence of adjustments and connections
If $\mathscr{G}$ is an $\mathsf{t}$-simply connected Lie groupoid over a compact and simply connected manifold $M$, then $\mathscr{G}$ admits a flat Cartan connection $\mathcal{H}$ if and only if $\mathscr{G}$ is isomorphic to an action Lie groupoid associated to a Lie group $\mathsf{H}$ acting on $M$.
@@
For more details on this point, see also the explicit discussion in [KS15b, near eqn. (9)].
@@ 7 Examples
We immediately note that Čech groupoids $\check{\mathscr{C}}(\sigma)$ for surjective submersions $\sigma:Y\rightarrow M$, cf. Example 2.1, are uninteresting: if $Y=\sqcup_{a}U_{a}$ for $\sqcup_{a}U_{a}$ an atlas of a manifold, the Lie algebroid of $\check{\mathscr{C}}(\sigma)$ is simply $Y$ itself, because the kernel of the map $\mathrm{d}\sigma:\mathrm{T}Y\rightarrow\mathrm{T}M$ is trivial in the fibers. Moreover, recall that $\check{\mathscr{C}}(\sigma)$ is Morita-equivalent to the groupoid $(M\rightrightarrows M)$, which produces a trivial special case of an action groupoid, see Section 7.3.434343We note, however, that Morita-equivalence of the gauge Lie groupoids does not have to imply physical equivalence, see e.g. [BJFJ+24, Sec. 4.4] for a counterexample in the context of higher gauge theory.
@@
However, as also argued in [Fis21a, Fis21b], the primitive may be not equivalent to a vanishing one, even locally.
@@ 7.1 Fundamental groupoid
$R_{\nabla}=-\nabla^{\mathrm{bas}}\zeta.$
@@
$\mathrm{d}^{\nabla^{\zeta}}\zeta=-\mathrm{d}^{\nabla^{\mathrm{bas}}}t_{\nabla^{\mathrm{bas}}}.$ (7.6)
@@ 7.2 Lie group bundles
There is a natural class of adjusted connections with non-trivial covariant adjustment. In order to understand this, we follow similar arguments as in the construction recipe I of [FLG24, see Ex. 1.12]: Assume an ordinary principal $\mathsf{H}$-bundle $P$, where $\mathsf{H}$ is a Lie group. Also assume that we have a manifold $\mathsf{T}$ on which $\mathsf{H}$ acts (from the left). Then we have two associated bundles:
@@
If $\mathsf{H}$ is semi-simple, then all covariant adjustments on $c_{\mathsf{H}}(P)$ are associated to a connection on $P$, and their primitive is the corresponding curvature on $P$.
Let us now assume that $\mathsf{T}=\mathbb{R}^{d}$, and that $\mathsf{H}$ acts faithfully on $\mathsf{T}$ with 0 as fixed point; this also implies that $\mathsf{H}$, and hence $c_{\mathsf{H}}(P)$, have a trivial center. Then there is a canonical singular foliation $\mathscr{F}$ on $\mathscr{T}$ with $M$ as a leaf generated by vector fields of the form
@@
Assume that the structural fiber $\mathsf{H}$ of a Lie group bundle is semi-simple. By Remark 7.9 and Lemma 7.11, there is a unique principal $\mathsf{H}$-bundle $P$ so that the Lie group bundle is of the form $c_{\mathsf{H}}(P)$, and its adjustments are the associated connections. By the previous discussion, $c_{\mathsf{H}}(P)$ is trivial if and only if $P$ is trivial. In this case, $P$ is flat, and its canonical flat connection induces a canonical flat connection on $c_{\mathsf{H}}(P)$ so that a field redefinition flattening a given adjustment always exists. Because of the non-trivial center, the only primitive is then $\zeta\equiv 0$. Thus, a semi-simple Lie group bundle is an ordinary setting (up to field redefinitions) if and only if the group bundle is trivial, if and only if $P$ is trivial.
@@ 7.3 Action groupoids
Recall that given a right action $(\mathbin{\vartriangleleft},\psi)$ of a groupoid $\mathscr{G}=(\mathsf{G}\rightrightarrows M)$ on a manifold $N$ (as in Definition 2.23 with $\mathscr{G}$ and $N$ concentrated in degrees 00 and, hence, with trivial differential), one can construct the corresponding action groupoid
with structure maps
for $g_{1,2}\in\mathscr{G}$ and $n\in N$ with $\psi(n)=\mathsf{t}(g)$ and $\mathsf{s}(g_{1})=\mathsf{t}(g_{2})$. Similarly, one can construct the action groupoid for left actions $(\vartriangleright,\psi)$, cf. [Mac05, Def. 1.6.10]. Infinitesimally, an action groupoid differentiates to an action Lie algebroid.
We now have the following statement, which allows us to construct natural connections on action groupoids.
@@
We assume that $\psi$ is a submersion so that $\mathfrak{X}(N)$ is generated by projectable vector fields. The proof thus canonically extends to all $\psi$ with this property on $\mathfrak{X}(N)$, for example, the proof also works for constant $\psi$. Assuming the existence of projectable generators can certainly be relaxed to a more general assumption: observe that the proof works for any smooth map $\psi$, if $\mathscr{G}$ is a Lie group bundle. That is, even more general, the proof works for $\psi$ admitting generators $\mathcal{V}$ of $\mathfrak{X}(N)$ such that $\left[\rho_{\mathfrak{k}}\left(\psi^{*}\nu\right),\mathcal{V}\right]$ are projectable vector fields for all $\nu\in\Gamma(\mathfrak{g})$.
@@
Let $P$ be the Hopf fibration $S^{7}\to S^{4}$, whose structural Lie group is given by $S^{3}\cong\mathrm{SU}(2)$, and define $\mathscr{T}\coloneqq\left(P\times\mathbb{R}^{3}\right)/S^{3}$. The canonical action algebroid $\mathscr{K}^{\rm t}$ inherited by the $c_{\mathsf{H}}(P)$-action on $\mathscr{T}$ admits an adjustment by Corollary 7.21 (given by a pullback of an adjustment on $c_{\mathsf{H}}(P)$), but there is no flat adjustment since $P$ would be otherwise trivial due to the fact that $S^{4}$ is simply connected. In particular, such adjustments are strict by Remark 7.22.
@@
A left section groupoid morphism $\chi\colon\mathscr{P}\mathsf{air}(M)\to\mathscr{G}$ of (7.21), also called a splitting, is in one-to-one correspondence to flat connections with trivial holonomy, inducing trivializations on $P$; $\chi$ is also called a (special type of) multiplicative Ehresmann connection on $\mathscr{G}\to\mathscr{P}\mathsf{air}(M)$. [^50] See [FM22] for a good introduction. If it exists, $\chi$ induces a $\mathscr{P}\mathsf{air}(M)$-action on $c_{\mathsf{H}}(P)$ simply given by the associated connection, a strict adjustment as discussed before, and now also the canonical flat connection (Maurer–Cartan form) w.r.t. the inherited trivialization of $c_{\mathsf{H}}(P)$. The splitting $\chi$ also induces a trivialization of $\mathscr{G}$ by
@@
where $\pi\colon\mathscr{G}\to\mathscr{P}\mathsf{air}(M)$ is the canonical projection induced by the fixed trivialization of $P$. This distribution is by construction a Cartan connection and admits a primitive $\zeta\coloneqq\pi^{!}\zeta^{M}$, the pullback of any primitive $\zeta^{M}$ of $\mathcal{H}^{M}$ as a form. Furthermore, there is always a strict covariant adjustment due to Remark 7.8. By a further pullback, making use of Corollary 7.20, we have a strict covariant adjustment on the action groupoid $\mathscr{K}$. ∎
@@ A Basics on graded manifolds
$\left.\bigodot\right.^{\bullet}V^{*}=~{}V^{0}~{}\oplus~{}V^{*}~{}\oplus~{}V^{*}\odot V^{*}~{}\oplus~{}\ldots~{},$ (A.2)
where $V^{0}$ denotes the ground field of the vector space and $\odot$ is the symmetric tensor product.
@@
The algebra of smooth functions on the graded manifold $\mathcal{X}=(X,\mathcal{O}_{X})$ is the algebra of smooth global sections of $\mathcal{O}_{X}$ and denoted by $C^{\infty}(\mathcal{X})$.
@@
is given by a pair $f=(|f|,f^{\sharp})$ consisting of a continuous map $|f|:X\rightarrow Y$ as well a ring homomorphism $f^{\sharp}:|f|^{*}\mathcal{O}_{Y}\rightarrow\mathcal{O}_{X}$, which is local in the sense that the stalks of $\mathcal{O}_{Y}$ at $f(x)$ are mapped to the stalks of $\mathcal{O}_{X}$ at $x$.
@@
$\mathsf{Hom}(\mathcal{X},\mathcal{Y})=\mathsf{Hom}(P(\mathcal{Y}),P(\mathcal{X}))=(\mathcal{Y}\otimes P(\mathcal{X}))^{0}~{},$ (A.5)
where $P(\mathcal{X})$ and $(-)^{0}$ denote the polynomial algebra of functions on $\mathcal{X}$ and the restriction to maps of degree 00, respectively. In the last equality, we used the fact that the contained map $f^{\sharp}$ in a morphism of graded manifolds $f=(|f|,f^{\sharp})$ is a ring homomorphism.
@@
In particular, in the category of sets, we simply have $\underline{\mathsf{hom}}(Y,Z)=\mathsf{Hom}(Y,Z)$. This also holds in the category of vector spaces, but it fails in the category of graded vector space. For example, let $V$, $U$, and $W$ be graded vector spaces with elements of positive degree. We have
@@
$\mathsf{Hom}(V\times U,W)=\mathsf{Hom}(V,\underline{\mathsf{hom}}(U,W))~{}~{}~{}\mbox{with}~{}~{}~{}\underline{\mathsf{hom}}(U,W)\cong W\otimes\left.\bigodot\right.^{\bullet}U~{}.$ (A.9)
@@
where we identified $C^{\infty}(\mathrm{T}[1]X)$ with the de Rham complex $\Omega^{\bullet}(X)$ on $X$ and $\mathrm{d}$ is the de Rham differential.

### v3

@@ 6.3 General statements on existence of adjustments and connections
If $\mathscr{G}$ is a $\mathsf{t}$-simply connected Lie groupoid over a compact and simply connected manifold $M$, then $\mathscr{G}$ admits a flat Cartan connection $\mathcal{H}$ if and only if $\mathscr{G}$ is isomorphic to an action Lie groupoid associated to a Lie group $\mathsf{H}$ acting on $M$.
@@
For more details on this point, see also the explicit discussion in [KS15b, near eqn. (9)]; see also [Bla13, Bla13] for further discussions.
@@ 7 Examples
We immediately note that Čech groupoids $\check{\mathscr{C}}(\sigma)$ for surjective submersions $\sigma:Y\rightarrow M$, cf. Example 2.1, are uninteresting: if $Y=\sqcup_{a}U_{a}$ for $\sqcup_{a}U_{a}$ an atlas of a manifold, the Lie algebroid of $\check{\mathscr{C}}(\sigma)$ is simply $Y$ itself, because the kernel of the map $\mathrm{d}\sigma:\mathrm{T}Y\rightarrow\mathrm{T}M$ is trivial in the fibers. Moreover, recall that $\check{\mathscr{C}}(\sigma)$ is Morita-equivalent to the groupoid $(M\rightrightarrows M)$, which produces a trivial special case of an action groupoid, see Section 7.3.474747We note, however, that Morita-equivalence of the gauge Lie groupoids does not have to imply physical equivalence, see e.g. [BJFJ+25, Sec. 4.4] for a counterexample in the context of higher gauge theory.
@@
However, as also argued in [Fis21a, Fis21b], the primitive may not be equivalent to a vanishing one, even locally.
@@ 7.1 Fundamental groupoid
$R_{\nabla}=-\nabla^{\mathrm{bas}}\zeta~{}.$
@@
$\mathrm{d}^{\nabla^{\zeta}}\zeta=-\mathrm{d}^{\nabla^{\mathrm{bas}}}t_{\nabla^{\mathrm{bas}}}~{}.$ (7.6)
@@ 7.2 Lie group bundles
There is a natural class of adjusted connections with non-trivial covariant adjustment. In order to understand this, we follow similar arguments as in the construction recipe I of [FLG24, see Ex. 1.12]: Assume an ordinary principal $\mathsf{H}$-bundle $P$, where $\mathsf{H}$ is a Lie group. Also assume that we have a manifold $\mathsf{T}$ on which $\mathsf{H}$ acts (from the left). [^49] For Section 7.2, $\mathsf{T}$ and its related constructions and assumptions are only needed to provide an extra geometric interpretation as in Table 1; from Example 7.12 to Example 7.15 one does not need this extra data. Then we have two associated bundles:
@@
If $\mathsf{H}$ is semi-simple, then all covariant adjustments on $c_{\mathsf{H}}(P)$ are associated to an Ehresmann connection on $P$, and their primitive is the corresponding curvature on $P$.
Semi-simplicity also implies that the correspondence in Lemma 7.11 is one-to-one. Slightly generalized, assuming that $\mathsf{H}$ has a trivial center implies that the associated connection on $c_{\mathsf{H}}(P)$ is in one-to-one correspondence with the Ehresmann connection on $P$. Thus, let us now assume that $\mathsf{H}$ and hence $c_{\mathsf{H}}(P)$ have a trivial center; furthermore assume that $\mathsf{T}=\mathbb{R}^{d}$, and that $\mathsf{H}$ acts faithfully on $\mathsf{T}$ with 0 as fixed point. Then there is a canonical singular foliation $\mathscr{F}$ on $\mathscr{T}$ with $M$ as a leaf generated by vector fields of the form
@@
Assume that the structural fiber $\mathsf{H}$ of a Lie group bundle is semi-simple. By Remark 7.9 and Lemma 7.11, there is a unique principal $\mathsf{H}$-bundle $P$ so that the Lie group bundle is of the form $c_{\mathsf{H}}(P)$, and its adjustments are the associated connections. In particular, $c_{\mathsf{H}}(P)$ is trivial if and only if $P$ is trivial. In this case, $P$ is flat, and its canonical flat connection induces a canonical flat connection on $c_{\mathsf{H}}(P)$ so that a field redefinition flattening a given adjustment always exists. Because of the non-trivial center, the only primitive is then $\zeta\equiv 0$. Thus, a semi-simple Lie group bundle is an ordinary setting (up to field redefinitions) if and only if the group bundle is trivial, if and only if $P$ is trivial.
@@ 7.3 Action groupoids
In the following we consider ordinary action groupoids, i.e. action groupoids in the sense of Definition 2.25, but with trivial differential and objects and morphisms concentrated in degrees 00.
We have the following statement, which allows us to construct natural connections on action groupoids.
@@
We assume that $\psi$ is a submersion so that $\mathfrak{X}(N)$ is generated by projectable vector fields. The proof thus canonically extends to all $\psi$ with this property on $\mathfrak{X}(N)$, for example, the proof also works for constant $\psi$. Thus, assuming the existence of projectable generators can certainly be relaxed to a more general assumption: observe that the existence of an action already implies that $\psi$ is a submersion by Eq. (7.14), if $\mathscr{G}$ is a transitive groupoid. Even more general, the proof works for $\psi$ admitting generators $\mathcal{V}$ of $\mathfrak{X}(N)$ such that
$\bigl{(}\nabla_{\mathfrak{k}}\bigr{)}^{\mathrm{bas}}_{\psi^{*}\nu}\left(\mathcal{V}\right)=\left(\left(\psi^{*}\nabla^{\rm{bas}}\right)_{\psi^{*}\nu}\bigl{(}\mathrm{d}\psi(\mathcal{V})\bigr{)}\right)^{\prime}$
for all $\nu\in\Gamma(\mathfrak{g})$, where one views $\mathrm{d}\psi(\mathcal{V})$ as a section of $\psi^{*}\mathrm{T}M$, and $\psi^{*}\nabla^{\rm bas}$ is the induced $\mathfrak{k}=\psi^{*}\mathfrak{g}$-connection on $\psi^{*}\mathrm{T}M$, given as the pullback of $\nabla^{\rm bas}$ as a $\mathfrak{g}$-connection on $\mathrm{T}M$ along the evident projection $\psi^{*}\mathfrak{g}\to\mathfrak{g}$; such a pullback exists here due to the fact that $\psi^{*}\mathfrak{g}\to\mathfrak{g}$ is a morphism of anchored vector bundles by Eq. (7.14). [^52] If the reader is unfamiliar with such pullbacks then see [Fis21b, Fis21c]. To do so, one only has to check whether $\left[\rho_{\mathfrak{k}}\left(\psi^{*}\nu\right),\mathcal{V}\right]$ projects “nicely” because the other summand in $\bigl{(}\nabla_{\mathfrak{k}}\bigr{)}^{\mathrm{bas}}_{\psi^{*}\nu}\left(\mathcal{V}\right)$ already behaves properly due to the fact that we define $\nabla_{\mathfrak{k}}$ as $\psi^{*}\nabla$.
The proof is left to the reader; in our corresponding examples $\psi$ will be a (surjective) submersion, so that we decided to keep the proof simpler.
@@
Let $P$ be the Hopf fibration $S^{7}\to S^{4}$, whose structural Lie group is given by $S^{3}\cong\mathrm{SU}(2)$, and define $\mathscr{T}\coloneqq\left(P\times\mathbb{C}^{2}\right)/S^{3}$. The canonical action algebroid $\mathscr{K}^{\rm t}$ inherited by the $c_{\mathsf{H}}(P)$-action on $\mathscr{T}$ admits an adjustment by Corollary 7.21 (given by a pullback of an adjustment on $c_{\mathsf{H}}(P)$), but there is no flat adjustment since $P$ would be otherwise trivial due to the fact that $S^{4}$ is simply connected. In particular, such adjustments are strict by Remark 7.22.
@@
A right section groupoid morphism $\chi\colon\mathscr{P}\mathsf{air}(M)\to\mathscr{G}$ of (7.20), also called a splitting, is in one-to-one correspondence to flat connections with trivial holonomy, inducing trivializations on $P$; $\chi$ is also called a (special type of) multiplicative Ehresmann connection on $\mathscr{G}\to\mathscr{P}\mathsf{air}(M)$. [^56] See [FM22] for a good introduction. If it exists, $\chi$ induces a $\mathscr{P}\mathsf{air}(M)$-action on $c_{\mathsf{H}}(P)$ simply given by the associated connection, a strict adjustment as discussed before, and now also the canonical flat connection (Maurer–Cartan form) w.r.t. the inherited trivialization of $c_{\mathsf{H}}(P)$. The splitting $\chi$ also induces a trivialization of $\mathscr{G}$ by
@@
where $\pi\colon\mathscr{G}\to\mathscr{P}\mathsf{air}(M)$ is the canonical projection induced by the fixed trivialization of $P$. This distribution is by construction a Cartan connection and admits a primitive $\zeta\coloneqq 0\oplus\zeta^{M}$, where 00 is the zero element of the Lie algebra of $\mathsf{H}$. Furthermore, there is always a strict covariant adjustment due to Remark 7.8. By a further pullback, making use of Corollary 7.20, we have a strict covariant adjustment on the action groupoid $\mathscr{K}$. ∎
@@ A Basics on graded manifolds
$\left.\bigodot\right.^{\bullet}V^{*}=~{}V^{0}~{}~{}\oplus~{}~{}V^{*}~{}~{}\oplus~{}~{}V^{*}\odot V^{*}~{}~{}\oplus~{}~{}\ldots~{},$ (A.2)
where $V^{0}$ denotes the ground field of the vector space and $\odot$ is the symmetric tensor product. The algebra of smooth functions $C^{\infty}(V)$ is obtained by the tensor product with $C^{\infty}(V_{0})$ over $P(V_{0})$.
@@
The algebra of smooth functions on the graded manifold $\mathcal{X}=(X,\mathcal{O}_{X})$ is the algebra of smooth global sections of $\mathcal{O}_{X}$ and denoted by $C^{\infty}(\mathcal{X})$. We will regard all our graded manifolds as smooth.
@@
is given by a pair $f=(|f|,f^{\sharp})$ consisting of a continuous map $|f|:X\rightarrow Y$ as well a ring homomorphism $f^{\sharp}:\mathcal{O}_{Y}\rightarrow|f|_{*}\mathcal{O}_{X}$, which is local in the sense that the stalks of $\mathcal{O}_{Y}$ at $f(x)$ are mapped to the stalks of $\mathcal{O}_{X}$ at $x$. Here, $|f_{*}|\mathcal{O}_{X}$ denotes the direct image sheaf.
@@
$\mathsf{Hom}(\mathcal{X},\mathcal{Y})=\mathsf{Hom}(C^{\infty}(\mathcal{Y}),C^{\infty}(\mathcal{X}))=(\mathcal{Y}\otimes C^{\infty}(\mathcal{X}))_{0}~{},$ (A.5)
where $C^{\infty}(\mathcal{X})$ and $(-)_{0}$ denote the smooth algebra of functions on $\mathcal{X}$ and the restriction to maps of degree 00, respectively. In the last equality, we used the fact that the contained map $f^{\sharp}$ in a morphism of graded manifolds $f=(|f|,f^{\sharp})$ is a ring homomorphism.
@@
In particular, in the category of sets, we simply have $\underline{\mathsf{hom}}(Y,Z)=\mathsf{Hom}(Y,Z)$. This also holds in the category of vector spaces, but it fails in the category of graded vector space. For example, let $V$, $U$, and $W$ be graded vector spaces concentrated in positive degrees. We have
@@
$\mathsf{Hom}(V\times U,W)=\mathsf{Hom}(V,\underline{\mathsf{hom}}(U,W))~{}~{}~{}\mbox{with}~{}~{}~{}\underline{\mathsf{hom}}(U,W)\cong W\otimes C^{\infty}(U)~{}.$ (A.9)
@@
where we identified $C^{\infty}(\mathrm{T}[1]X)$ with the de Rham complex $\Omega^{\bullet}(X)$ on $X$ and $\mathrm{d}$ is the de Rham differential, cf. Definition 2.16.